# Video[Video][video]

### v2

@@ 3.6 Summary of transfer functions
4 Multi-contrast multi-channel reconstructions across length scales
4.1 Label-free simulated and real test specimen
We tested our newly developed anisotropic label-free model on polarization-diverse data acquired from a laser-etched anisotropy phantom with transverse radially anisotropic bubbles arranged in a spoke pattern (Figure 3a,i). We measured four volumetric Stokes datasets (Figure 3a,ii) and applied waveOrder’s reconstruction algorithm to estimate three label-free material properties that correspond to phase and transverse birefringence (Figure 3a,iii).
In parallel, we simulated the anisotropic phantom (Figure 3b,i) and the image formation process (Figure 3b,ii), then we applied an identical reconstruction algorithm to estimate material properties (Figure 3b,iii). Figures 3a, ii-ii and b,ii–iii can be directly compared to indicate the quality of our models, where differences can arise from imperfect modeling of both the object and the image formation process. While our simulations recreate the most important contrast features, the real measurements have contrast with a broader axial extent and poorer transverse spatial resolution than our simulations—likely due to imperfections in our phantom and slightly aberrated imaging.
We compared an earlier ray-optics based voxel-by-voxel reconstruction algorithm [3] with waveOrder’s new wave optical reconstruction algorithm Figure 3c. We find that wave-optical reconstructions yield marginally improved transverse resolution Figure 3d measured via transverse modulation transfer functions from azimuthal profiles, and denoised and defocus-symmetric axial profiles Figure 3e. We also observe orientation reversals between spokes, reconstruction artifacts that are analogous to well-known negative ringing artifacts in fluorescence deconvolution.
@@ 4.1 Label-free simulated and real test specimen
Figure 3: Demonstration of vectorial multi-channel reconstruction with experiment and simulation. We imaged (a) (i) a laser-etched spoke pattern of transverse radially anisotropic bubbles. (ii) We made volumetric measurements of the Stokes parameters (columns), and applied a multi-channel label-free reconstruction to (iii) recover material properties. (b) We (i) simulated the phantom’s material properties, (ii) simulated the imaging process, then (iii) simulated the reconstruction. (c) We encoded transverse birefringence properties into color, where brightness indicates the strength of the anisotropy and hue indicate the slow-axis orientation. We compared the orientation and phase reconstructions (rows) in ray- and wave-optics reconstructions (columns). (d) For each data and material property, we measured a series of azimuthal profiles at different radii on the spoke pattern and used the 10th–90th percentile modulation as an empirical estimate of the transverse modulation transfer function (MTF). (e) Similarly, we measured axial profiles through each data and material property. See also Video 3.
4.2 Cells and tissues
We applied waveOrder reconstructions to multi-channel data acquired from samples across length scales. Figure 4 shows alternating columns of data and reconstructions for transverse birefringence, phase, and fluorescence density. In data acquired from A549 cells (Figure 4a) we observe improved sectioning, denoising, and contrast in phase and reconstructed fluorescence properties compared to their raw-data counterparts (Figure 4a, ii–iii). In the orientation channel (Figure 4a, i) we observe marginal improvements in contrast, but generally poor performance with reduced SNR and suppression of features that are apparent in the raw data. We attribute some of the performance drop to our imperfect noise model—we reconstruct from non-Gaussian Stokes parameters which is at odds with our Tikhonov least-squares reconstruction algorithm. Additionally, we have not explored the interaction between our Stokes-based background correction and our wave-optical reconstructions, another likely area for improvement.
We acquired multi-contrast data from an entire living zebrafish (Figure 4b), then reconstructed fluorescence density, phase, and birefringence from specific regions of interest (Figure 4c–e). We observe improved sectioning, denoising, and contrast in all three reconstructions. For example, in the label-free Figure 4i, c–e we see improved contrast between the gut and muscles, and in the fluorescence channels Figure 4iii, c–e we see improved contrast and resolution of immune cells. Improved contrast in the zebrafish gut (bottom of Figure 4c) is particularly valuable for tracking immune-cell dynamics encoded in the fluorescence reconstructions.
Figure 4: Reconstructions improve visualization and interpretation across length scales in (a) A549 cells, (b) zebrafish, and (c–e) zebrafish regions of interest. (i) Orientation data (left) and reconstructions (right). (ii) Brightfield data (left) and phase reconstructions (right). (iii) Fluorescence data (left) and reconstructions (right). See also Videos 4 and 5.
4.3 Choice of regularization parameter
waveOrder’s physics-informed reconstructions rely on a list of fixed parameters that specify imaging conditions (e.g., numerical apertures, wavelength, voxel sizes) and a single regularization parameter $\eta$, which penalizes large solutions and suppresses noise. Though [35, 36], in practice, it is common to inspect a regularization sweep and choose an empirical solution (for example, Figure 5 and Figure S1). Regularization parameters that are too small result in noisy reconstructions, while regularization parameters that are too large result in blurry reconstructions. Additionally, we find that the transverse mid-band power, the power in transverse spatial frequencies between one-eighth and one-quarter of the cutoff, is a useful scalar metric that can indicate high-quality reconstructions—see the plateau in Figure 5(d).
waveOrder’s current design uses a single regularization parameter per contrast mode. For example, we use one regularization parameter for each fluorescence deconvolution, one for each phase-from-brightfield reconstruction, and one for each joint reconstruction of phase and orientation. These single-parameter Tikhonov-regularized least squares reconstructions assume that all measured channels have an SNR proportional to their intensity and that the physical model accurately describes the differences in intensity between channels. This requirement becomes challenging when reconstructing directly from Stokes parameters, typically estimated with varying SNR. For example, the Stokes parameters that preceded the reconstruction in Figure 5 include a very noisy $S_{3}$. Here, we dropped $S_{3}$ from our dataset, ignoring its small contribution, but improved reconstructions with channel-specific regularization parameters may allow us to use this channel more effectively.
In future work, we are excited to pursue data-adaptive and channel-adaptive regularization in a data- and task-dependent manner. We are especially excited by methods that can address model mismatch, which often limits the quality of our reconstructions. For example, while Figure 5(iii,b–e) shows improved contrast and sectioning, model mismatch limits the quality of our reconstructions by leaving artifactual double cones throughout the reconstruction, particularly visible as transverse rings from bright out-of-focus cells.
Figure 5: Effect of regularization parameter on reconstructions. (a) Zebrafish neuromasts imaged in fluorescence and brightfield channels are (b) reconstructed over a range of regularization parameters, with small parameters generating noisy results and large parameters generating blurred results. Near the empirically selected optimal reconstruction (b, ii), the DC power (c) is nearly constant. In contrast, the mid-band power (d), the power for transverse spatial frequencies between one-eighth and one-quarter of the cutoff, changes when the regularization parameter is too high or too low and plateaus when the regularization parameter is in the correct range.
5 Limitations
While waveOrder enables the development of image formation models and reconstruction algorithms for a wide variety of computational microscopy techniques, its current design makes the following assumptions that are not valid in certain applications:
  • consistent SNR in all channels, which assumes that all data channels that encode a given contrast have a similar signal-to-noise ratio.
  • channel linearity, which excludes non-linear crosstalk between channels e.g. FRET, strongly scattering samples,
  • spatial linearity, which excludes saturation of fluorescent samples or the camera,
  • shift invariance, which excludes shift-variant point spread functions [37],
  • weak, single scattering, which excludes thick multiply scattering samples [38], e.g. older zebrafish and thick tissue slices,
  • aberration-free imaging, which excludes applications in the presence of sample-induced aberrations.
  • contrast-separable imaging, which excludes crosstalk between label-free and fluorescence contrast modes [39].
6 Discussion and conclusion
waveOrder improves on earlier multi-channel correlative imaging methods in several ways. Compared to LC-PolScope reconstruction methods [40], waveOrder uses a wave-optical approach that reconstructs from diffraction-limited groups of voxels instead of reconstructing voxel by voxel. waveOrder also extends quantitative label-free imaging with phase and polarization (QLIPP) [3], by using a vectorial wave-optics model of multiple specimen properties, not just phase. waveOrder also extends permittivity tensor imaging (PTI) [21], which reconstructs only uniaxial permittivity tensors while waveOrder is compatible with arbitrary biaxial materials. The waveOrder framework also generalizes previous work on imaging polarized fluorescence ensembles [14, 24], where diffraction-limited ensembles of fluorescent dipoles reduce to orientation distribution functions. As discussed earlier, we view careful noise handling and more robust background correction as areas for future work.
To conclude, we find that linear models provide a strong framework for widely applicable computational microscopy techniques, including phase, absoprtion, birefringence, diattenuation, and anistropic fluorescence imaging. We find the waveOrder framework useful for understanding, simulating, and reconstructing data acquired with this class of techniques, and we demonstrate its ability to improve multi-contrast multi-channel data across length scales.
7 Data and code availability
Demos, data and code can be found in the recOrder and
waveOrder repositories.

### v3

@@ 4.2.8 Summary of transfer functions
4.2.9 Reconstruction algorithm
For each contrast mode, the forward operators in Eq. 10 can be decomposed using the singular value decomposition
$\mathcal{H}=\sum_{i}^{R}s_{i}\mathbf{u}_{i}\mathbf{v}^{\dagger}_{i},$ (15)
where $R$ is the rank of $\mathcal{H}$, $\{\mathbf{u}_{i}\}$ is a set of orthonormal data-space vectors that span the space of deterministic expected data, $\{\mathbf{v}_{i}\}$ is a set of orthonormal object-space vectors that span the measurement space of object properties, and $\{s_{i}\}$ are singular values.
We use this decomposition to solve Eq. 12 in a single step for each mode
$\mathcal{R}_{\eta}=\sum_{i}^{R}\frac{s_{i}}{s_{i}^{2}+\eta}\mathbf{v}_{i}\mathbf{u}^{\dagger}_{i}.$ (16)
In practice we choose a different regularization parameter $\eta$ for each contrast mode.
4.3 Cell-scale optical pooled screen data
Auto-tuned reconstructions 2D phase reconstructions were computed from multi-plane volumetric data using a phase-from-defocus forward model. For each field of view, the volumetric data were divided into a $6\times 6$ grid of tiles with 20% overlap.
To optimize reconstruction quality, we implemented gradient-based auto-tuning of the model parameters. The objective function maximized the power spectral density in a mid-band spatial frequency range (1/8–1/4 of the transverse cutoff frequency), which corresponds to cellular feature sizes while avoiding low-frequency background variations and high-frequency noise. Three parameters were optimized: axial focal offset ($z_{\text{offset}}$), illumination zenith tilt angle ($\theta_{\text{zenith}}$), and azimuthal tilt angle ($\phi_{\text{azimuth}}$).
The optimization used the Adam algorithm [kingma_adam_2014] with learning rates of 0.02 µm, 0.005 rad, and 0.001 rad per iteration for $z_{\text{offset}}$, $\theta_{\text{zenith}}$, and $\phi_{\text{azimuth}}$, respectively. Each tile was optimized for 100 iterations, with the reconstruction and gradient computation performed in PyTorch [paszke_pytorch_2019] to enable automatic differentiation. The regularization strength for the inverse problem was fixed at $\eta=10^{-2}$. Fixed optical parameters included illumination wavelength (450 nm), detection NA (0.15), illumination NA (0.1), and sample refractive index (1.0 for air-mounted samples). Optimization progress was monitored using TensorBoard, logging parameter values, loss curves, and reconstructed phase images at each iteration.
Cell segmentation Cell segmentation was performed on reconstructed phase images to quantitatively compare reconstruction methods. Three conditions were evaluated: (i) raw multi-plane intensity data from the highest-contrast slice $z$ plane, (ii) nominal reconstruction with zero illumination tilt, and (iii) auto-tuned reconstruction with optimized illumination parameters. All images were cropped to a standardized $300\times 300$ pixel region for analysis.
Segmentation was performed using Cellpose cyto2 model [stringer_cellpose_2021, pachitariu_cellpose_2022]. Cell diameters were set to 175 pixels across all conditions and fields of view, with a flow threshold of 0.9. Ground truth annotations were manually created for four representative image regions and saved as labeled arrays.
Segmentation accuracy was quantified using intersection-over-union (IoU) metrics with Hungarian matching. For each image pair (predicted and ground truth), we computed the IoU between all object pairs and performed optimal one-to-one matching using linear sum assignment to maximize total IoU. This yielded a distribution of per-object IoU scores for each condition.
We computed recall (fraction of ground truth cells successfully detected) and precision (fraction of predicted cells matching ground truth) as functions of IoU threshold $t\in[0,1]$, sampled at 0.01 intervals. F1 scores were calculated as $\text{F1}(t)=2\cdot\text{precision}(t)\cdot\text{recall}(t)/(\text{precision}(t)+\text{recall}(t))$ at each threshold. Precision-recall-F1 curves were generated for each field of view to characterize segmentation performance across the three reconstruction conditions.
4.4 Organelle-scale optical pooled screen data
Tiling and stitching: Each field of view was divided into a $5\times 5$ grid of tiles with 25 pixels of overlap. After reconstruction, the tiles were stitched with linearly weighted Euclidean distance transform (EDT)-based blending.
Auto-tuned reconstructions: Raw brightfield through-focus stacks were processed using a scalar reconstruction routine to generate 3D phase density maps. Additionally, a phase-from-defocus model was auto-tuned by optimizing the axial focal offset ($z_{\text{offset}}$) for power spectral density in a mid-band spatial frequency range (1/8 –1/4 of the transverse cutoff frequency), summarizing the cellular phenotype into a single high-SNR plane.
Image registration: To account for chromatic aberrations and mechanical shifts between fluorescence (ground truth) and label-free channels, we implemented an automated registration routine. For every field of view, a translational offset sweep ($\pm 5$ pixels in $X$ and $Y$) was performed. The optimal offset was determined by maximizing the spatial overlap between the Frangi-filtered binary masks of the fluorescence channel and the label-free reconstruction. This offset was applied prior to feature extraction to ensure accurate spatial comparison.
Organelle segmentation: Organelles were segmented using a multi-scale Frangi vesselness filter adapted for distinct morphologies [lefebvre_nellie_2025]. Prior to filtering, label-free images (raw, 3D reconstruction, 2D reconstruction) were pre-processed with Contrast Limited Adaptive Histogram Equalization (CLAHE; clip limit = 0.01, kernel size = $64\times 64$ pixels) to normalize local contrast.
Mitochondria: To capture tubular networks, we utilized a Frangi filter with high plate-sensitivity ($\alpha=4.0$, $\beta=0.5$) across scales corresponding to radii of 0.10.11.51.5 µm.
Endosomes: To capture vesicular/blob-like structures, we utilized a filter with isotropic sensitivity ($\alpha=0.5$, $\beta=0.5$) across radii of 0.20.21.51.5 µm.
Probability maps were thresholded using a combined Triangle/Otsu method on the logarithmic vesselness response [lefebvre_nellie_2025]. A cytoplasm mask was generated by eroding the cell segmentation mask (2 iterations) and excluding the nucleus (dilated by 20 pixels) to restrict analysis to the cytoplasm and remove nuclear artifacts.
Quantitative comparison: Segmentation performance was quantified using a spatial overlap metric, defined as the percentage of the binarized fluorescent ground truth area that intersected with the label-free segmentation mask.
Statistical analysis: Statistical comparisons between imaging modalities (raw, 3D reconstruction, 2D reconstruction) were performed treating individual cells as biological replicates. We used a repeated measures ANOVA to test for differences across reconstruction methods, followed by post-hoc paired $t$-tests with Bonferroni correction for pairwise comparisons.
4.5 Neuromast segmentation and classification
Segmentation and annotation: Neuromast nuclei were manually segmented using signal from the she:H2B-EGFP transgenic line [peloggia_adaptive_2021]. Each nucleus in a neuromast was assigned a unique identifier. Nuclei were assigned to one of two categories. The first category, mantle cells, form the boundary of each neuromast. The second category, hair and support cells, contains the internal proliferative and mechanosensory cells of the neuromast.
Texture metric calculation: Single cell Shannon Entropy values were calculated using the neuromast nuclear instance segmentations described above [shannon_mathematical_1948]. Similarly, homogeneity and correlation values were calculated to have second order metrics describing fluorescent signal texture(s) [haralick_textural_1973].
Statistical analysis: Kruskal-Wallis H-tests for independent samples were computed to compare first and second order texture measurements between imaging modalities and reconstructions [kruskal_use_1952], with significance levels (*) $p<0.05$, (**) $p<0.01$, (***) $p<0.001$.
4.6 Cardiomyocyte reconstructions and profiles
Cardiomyocytes were imaged with polarization-resolved label-free microscopy with 8 sector-shaped illumination patterns and a single full-aperture illumination pattern [yeh_permittivity_2024]. Raw intensity measurements acquired with four polarization states were averaged to reduce polarization-dependent artifacts, then reconstructed into 3D phase volumes. Two reconstruction strategies were compared: phase reconstruction from a circular illumination aperture versus phase reconstruction combining all nine illumination apertures. Manually drawn line profiles across cellular structures were extracted to quantitatively compare image quality between reconstruction approaches.
4.7 Zebrafish embryo analysis
Preprocessing: Raw volumes were deskewed and registered by estimating affine transformations for each view to correct the oblique geometry and align all views into a common coordinate system.
Tail: Straightening and profiles: The zebrafish tail was straightened into rectilinear coordinates using curvilinear resampling. A manually-annotated midline was smoothed with a cubic spline and parameterized by arc length $s$. At each position along the midline, we computed the unit tangent vector $\mathbf{t}(s)$ and perpendicular normal vector $\mathbf{n}(s)$. For the label-free and nuclei channels, we sampled symmetrically in both directions (dorsal and ventral) from the midline. For membrane fluorescence, we sampled only the dorsal side to isolate the notochord region. Local perpendicular widths were computed adaptively by ray-casting along normals until reaching the mask boundary, and the straightened coordinate system was normalized such that $n\in[0,1]$ spans from midline to dorsal edge. Image intensities were resampled onto a regular grid in curvilinear $(s,n)$ space using bilinear interpolation.
Spatial distribution quantification: The straightened image was divided into anterior and posterior halves along the midline axis. For each half, we computed the total integrated intensity by summing all pixels. The spatial enrichment ratio was calculated as $R=I_{\text{anterior}}/I_{\text{posterior}}$.
Statistics: Standard errors for the enrichment ratio were estimated via bootstrap resampling (n=10,000n=10{,}000 iterations). For each bootstrap sample, we resampled rows (perpendicular to the midline, i.e., along the dorsal axis) with replacement, recomputed the integrated intensities for each half, and calculated the ratio. The standard error of the mean was computed as SEM = $\sigma_{\text{boot}}/\sqrt{n}$, where $\sigma_{\text{boot}}$ is the standard deviation of the bootstrap distribution.
Notochord: ROI selection and extraction: Three sub-ROIs (posterior, median, and anterior) were manually selected from the tail region, and rotated rectangular regions were resampled onto a regular grid using bilinear interpolation.
Profile extraction: Intensity profiles were extracted along the long axis (horizontal center) of each straightened sub-ROI. Profiles were computed over the central region spanning 10–40% of the ROI length, averaging intensities across 5 pixels perpendicular to the profile direction. For each profile line, the sum of intensities across the 5-pixel width was computed, yielding a 1D profile representing integrated signal along the notochord axis.
Peak detection and spacing quantification: Peaks in the phase reconstruction profiles were manually identified from the normalized intensity traces, and distances between consecutive peaks were computed.
Statistics: For each region (posterior, median, anterior), mean peak spacing was calculated with standard error of the mean (SEM = $\sigma/\sqrt{n}$) and standard deviation (SD). Statistical comparisons between regions were performed using the Mann-Whitney U test (two-sided) with significance levels (*) $p<0.05$, (**) $p<0.01$, (***) $p<0.001$.
Eye: Annotation and resampling: Similar to the tail, we manually annotated the retina boundary and midline, resampled onto a Cartesian grid, then plotted 1D projections along anatomical axes.
Profiles: The resampled retina was divided into basal, medial, and apical thirds. For each zone and fluorescence channel, we computed total integrated intensity $I_{\text{zone}}$ and normalized to the medial zone: $I_{\text{zone}}^{\text{rel}}=I_{\text{zone}}/I_{\text{medial}}$.
Statistics: Standard errors were estimated by computing per-column mean intensities within each zone, normalizing to the medial intensity, and calculating SEM = $\sigma/\sqrt{M}$ across $M$ valid columns.
5 Extended data
Method Name Contrast Optics Model Lang. Autodiff. Library Object Rep. Recon. Method Learning Architecture Imaging Types Blind Shift Variant Ref
WaveOrder Fl, VFl, Ph, Bir, Abs Part. coh. vect., single scattering Python PyTorch Voxel Tk PI-FC WF, LS, DPC, OPM, Conf, Pol. oblique ill. Yes Yes This work
[…]
AOViFT Fl Scalar Python Tensorflow Voxel RL ViT LLS, WF Yes Yes [alshaabi_fourier-based_2025]
[…]
Deconwolf Fl Scalar C None Voxel RL+ None WF No No [wernersson_deconwolf_2024]
[…]
RDM Fl Scalar Python PyTorch Voxel LS U-Net WF, LS, miniscope, endoscope No Yes [kohli_ring_2025]
[…]
PTI Ph, Bir Part. coh. vect. Python None Voxel Tk None Pol. oblique ill. No No [yeh_permittivity_2024]
[…]
DeepInverse General General Python PyTorch General General General MRI, CT, optics, general Yes Yes [tachella_deepinverse_2025]
Extended Data Table 1: Comparison of computational imaging and reconstruction frameworks. Abbreviations: Fl = fluorescence; VFl = vector fluorescence; Ph = phase; Bir = birefringence; Abs = absorption; Part. coh. vect. = partially coherent vectorial; Tk = Tikhonov; RL = Richardson–Lucy; RL+ = accelerated RL; RLN = Richardson–Lucy network; LS = least squares; FC = fully connected; PI = physics-informed; PI-CNN = physics-informed CNN; INR = implicit neural representation; MLP = multi-layer perceptron; ViT = vision transformer; WF = widefield; LS (imaging) = light sheet; DPC = differential phase contrast; OPM = oblique-plane microscopy; Conf = confocal; Pol. obl. ill. = polarization-resolved oblique illumination; LLS = lattice light sheet; SIM = structured illumination microscopy; iSIM = instant SIM; diSPIM = dual-view inverted selective plane microscopy; STED = stimulated emission depletion; holo. = holography; ptych. = ptychography.
Extended Data Figure 1: Physically interpretable bases for modeling material properties and detected light. (a) Coherent scatterers are described by a complex-valued scattering potential tensor that encodes how an incident field induces dipole moments within the specimen. Expanding this tensor into its spherical-harmonic components separates isotropic, circularly anisotropic, and linearly anisotropic responses (rows). The real parts of the expansion coefficients correspond to phase and birefringence, and the imaginary parts correspond to absorption and diattenuation. (b) Incoherent (fluorescent) scatterers are represented by emission dipoles oriented along $\mathbf{\hat{x}}$, $\mathbf{\hat{y}}$, or $\mathbf{\hat{z}}$. (c) Detected light is represented by Stokes parameters, allowing birefringence and diattenuation to be modeled within the same operator framework. The imaging operator linking (a–b) material properties to (c) measurable intensities is denoted $\mathcal{H}$, and its pseudo-inverse is denoted $\mathcal{H}^{+}$.
Extended Data Figure 2: Key components of the forward model. (a) A dipole emitter (blue arrow) emits polarized fields (red arrows) in an anisotropic pattern (black line, radius proportional to power). (b) Arbitrary linear, circular, and elliptical dipole emissions can be modeled with the Green’s tensor spectrum $G_{ij}(\boldsymbol{\nu})$. Each surface shows the quarter-maximum intensity of on-shell spectral field components (rows) $e_{x}$, $e_{y}$, and $e_{z}$ emitted by dipoles oriented along (columns) $\hat{\textbf{x}}$, $\hat{\textbf{y}}$, and $\hat{\textbf{z}}$, with relative phase shown in color (see color rose). Orthogonal slices through the Green’s tensor spectrum and the Green’s tensor are shown in Supp. Fig. 2c–g. 3D support of various transfer functions with phase encoded in color. (c) Scalar label-free transfer function with $\text{NA}_{\text{ill}}=0.5$ and $\text{NA}_{\text{det}}=0.75$. (d) Scalar fluorescence transfer function with $\text{NA}_{\text{det}}=0.75$. (e–f) Label-free vector transfer functions with $\text{NA}_{\text{det}}=0.75$ and circularly polarized illumination that expresses polarization-resolved data (rows) as outputs of filters that modulate material properties (columns) with (e) low-NA illumination $\text{NA}_{\text{ill}}=0.1$ and (f) high-NA illumination $\text{NA}_{\text{ill}}=0.5$. (g) Vectorial fluorescence transfer functions with $\text{NA}_{\text{det}}=0.75$ and circularly polarized illumination that expresses polarization-resolved data (rows) as outputs that modulate material properties (columns), here spherical harmonic components of fluorescent dipole orientation distribution functions.
Extended Data Figure 3: Network architecture for auto-tuned phase-from-defocus reconstructions. (a) Illumination-parameter branch: illumination numerical aperture, tilt zenith, and tilt azimuth (3 parameters) are used to generate the illumination pupil ($\mathtt{Y,X}$) via calc_ill_pupil. (b) Detection-parameter branch: detection numerical aperture, defocus offset, and detection Zernike coefficients (2 + N_z) parameters are used to calculate the detection pupil (Z,Y,X) through calc_det_pupil. (c) Transfer-function construction: illumination and detection pupils are combined using calc_tf to form a 3D defocus transfer function (Z,Y,X). (d) Inverse imaging model: the transfer function and raw defocused intensity stack (Z,Y,X) are passed to apply_inverse_tf, producing the phase reconstruction (Y,X) with a fixed regularization parameter. (e) Mid-band power loss: reconstructed phase is evaluated by calc_midband_power to compute the scalar loss for optimization. Arrows denote data flow; dashed arrows denote backpropagation to optimize the input parameters; dimensionalities are below each data object. This architecture jointly optimizes the phase reconstruction, illumination, and detection parameters from raw defocused measurements.
Extended Data Figure 4: Quantitative evaluation of auto-tuned reconstructions on optical pooled screen data. (a) Raw label-free images from representative tiles across the 35 mm well show variation in illumination and contrast between center and edge tiles. (b) Manual segmentations used as ground truth. (c) CellPose segmentations on raw data. (d) Segmentations on nominal (untuned) reconstructions. (e) Segmentations on physics-informed, auto-tuned reconstructions. (f–g) Quantitative evaluation of segmentation accuracy, reported as recall and precision scores versus IoU threshold for raw (red), nominal (blue), and auto-tuned (green) reconstructions.
Extended Data Figure 5: Label-free light path schematics. (a) Straight path: light from the source (S) is focused by the condenser (C) and transmitted through the sample, which is moved along the $z$-axis. Transmitted light is collected by the first objective (O1) and imaged with a tube lens (TL) onto a camera. (b) Oblique detection path: the illumination remains identical to (a), but detection is re-imaged at an oblique angle using a secondary objective (O2) and tertiary objective (O3). (c) Oblique illumination path: the detection path remains straight as in (a), but illumination is provided obliquely by masked illumination sectors (yellow) in the source plane. S = source; C = condenser; O = objective; TL = tube lens. Blue dots indicate the sample plane; $z$ indicates the sample translation axis.
Extended Data Figure 6: Quantitative phase reconstruction enhances label-free organelle segmentation. (a) Mitochondria segmentation in live A549 cells. (i) Comparison of imaging modalities: ground-truth fluorescence (mCherry–Mito), raw brightfield (center slice), 3D phase reconstruction (center slice), and 2D phase reconstruction (auto-tuned). Rows show raw intensity data, segmentation overlays (cyan), color-coded connectivity labels, and overlays of segmentation (yellow) between ground-truth fluorescence (red) and reconstructions (green). (ii) Zoomed regions highlight the recovery of fine mitochondrial networks in phase reconstructions relative to raw brightfield. (iii) Quantification of segmentation overlap, reported as the percentage of the total fluorescent area covered by the label-free segmentation ($N=11$ cells). 2D phase reconstruction (38% $\pm$ 5) significantly outperforms raw brightfield (9.4% $\pm$ 4). (b) Endosome segmentation in A549 cells. (i) Comparison of imaging modalities as in (a) using GFP-labeled endosomes. (ii) Zoomed regions show improved contrast of vesicular structures in raw and phase-reconstructed images. (iii) Quantification of segmentation overlap shows similar improvements for endosomes ($N=11$ cells). Statistical analysis used repeated-measures ANOVA with post-hoc paired $t$-tests and Bonferroni correction. Box plots indicate mean $\pm$ s.d.; significance levels: (*) $p<0.05$, (**) $p<0.01$, (***) $p<0.001$. Scale bars: 20 µm.
Extended Data Figure 7: Neuromast cell classification. (a)(i) Shannon Entropy % change during reconstruction of oblique label-free detection (top left), oblique fluorescence detection (top right), straight label-free (bottom left) and straight fluorescence (bottom right). (a)(ii) Violin plot showing the distributions of single cell Shannon Entropy values for straight and oblique detection both before and after reconstruction. (b)(i) Correlation % change for oblique (top) and straight (bottom) detection during reconstruction. (b)(ii) Violin plot showing the distributions of single cell correlation values for mantle (green) and hair/support (black) cells imaged with straight or oblique detection both before and after reconstruction.
Extended Data Figure 8: Multispectral iPSC data. (a) Configuration for multispectral imaging of iPSCs and iPSC-derived neurons (iNeurons), combining a label-free channel with multiple fluorescence channels. Linear unmixing enables organelle-specific interpretation. (b) Representative label-free images (left) and reconstructed phase. (c) Restored and unmixed images (columns) across differentiation stages (rows), demonstrate a rich mapping between the dense label-free channel and the unmixed fluorescence channels. (d) Comparing brightfield (far left) and reconstructed phase (far right) to three fluorescence organelle channels (center) with overlays (center left and right) show improved correlation between reconstructed phase and fluorescence, particularly in the lipid droplet channel (green arrows).
@@ 5 Extended data
Extended Data Figure 9: Demonstration of vectorial multi-channel reconstruction with experiment and simulation. We imaged (a) (i) a laser-etched spoke pattern of transverse radially anisotropic bubbles. (ii) We made volumetric measurements of the Stokes parameters (columns), and applied a multi-channel label-free reconstruction to (iii) recover material properties. (b) We (i) simulated the phantom’s material properties, (ii) simulated the imaging process, then (iii) simulated the reconstruction. (c) We encoded transverse birefringence properties into color, where brightness indicates the strength of the anisotropy and hue indicate the slow-axis orientation. We compared the orientation and phase reconstructions (rows) in ray- and wave-optics reconstructions (columns). (d) For each data and material property, we measured a series of azimuthal profiles at different radii on the spoke pattern and used the 10th–90th percentile modulation as an empirical estimate of the transverse modulation transfer function (MTF). (e) Similarly, we measured axial profiles through each data and material property. See also Video 5.
[…]
newinmethods segment=2 and not segment=1
6 Biological materials availability
All unique biological materials classified are available from authors upon reasonable request or from commercial sources.
7 Data availability
Demo data with example reconstructions are available in the WaveOrder repository.
8 Code availability
Code can be found in the WaveOrder repository.